# Adaptive quantum dynamics with the time-dependent variational Monte Carlo method

### v3

@@ Abstract
We introduce an extension of the time-dependent variational Monte Carlo (tVMC) method that adaptively controls the expressivity of the variational quantum state during the simulation. This adaptive tVMC approach addresses numerical instabilities that arise when the variational ansatz is overparameterized or contains redundant degrees of freedom. Building on the concept of the local-in-time error (LITE)—a measure of the deviation between variational and exact evolution—we introduce a procedure to quantify each parameter’s contribution to reducing the LITE, using only quantities already computed in standard tVMC simulations. These relevance estimates guide the selective evolution of only the most significant parameters at each time step, while maintaining a prescribed level of accuracy. We benchmark the algorithm on quantum quenches in the one-dimensional transverse-field Ising model using both spin-Jastrow and restricted Boltzmann machine wave functions, with an emphasis on overparameterized regimes. The adaptive scheme significantly improves numerical stability and reduces the need for strong regularization, enabling reliable simulations with highly expressive variational ansätze.
@@ I Introduction
Understanding the dynamics of strongly correlated quantum many-body systems remains a fundamental and long-standing challenge in physics. In recent years, this problem has attracted renewed interest, driven by theoretical and experimental advances in the study of out-of-equilibrium phenomena [1, 2, 3, 4, 5, 6, 7, 8], as well as by the rapid development of quantum technologies [9, 10, 11]. In particular, accurately describing quantum dynamics is essential to understanding the propagation of entanglement and coherence, and ultimately, to achieving control over them [12]. Exact solutions to the time evolution of quantum many-body systems are limited to a few analytically tractable or numerically trivial cases. The primary obstacle is the exponential growth of the Hilbert space with the number of degrees of freedom, which makes a full description of the wave function computationally intractable. Consequently, a variety of numerical methods have been developed to circumvent this limitation, including tensor network techniques, quantum Monte Carlo, and variational approaches based on neural-network quantum states, all of which aim to efficiently parametrize the wave function in regimes where exact methods become intractable [13, 14, 15, 16, 17, 18].
@@
In this work, we focus on the time-dependent variational Monte Carlo (tVMC) method [19, 20, 21], which approximates the system’s wave function using a variational ansatz and, as described below, evolves it according to a variational principle. A key feature of tVMC is its flexibility in treating systems of arbitrary dimensionality, as its formulation does not rely on a specific geometry and connectivity. Moreover, since the wave function is represented by a parametrized variational ansatz, the computational cost of the method depends on the number of variational parameters rather than on the dimension of the Hilbert space, allowing for favorable scaling even in large or high-dimensional systems. This makes it complementary to other powerful methods, such as tensor networks [22, 13, 14, 15, 23, 24, 25], which are most effective in low-dimensional systems or when the entanglement is sufficiently constrained. However, the accuracy of a tVMC simulation crucially depends on the expressivity of the chosen variational ansatz. Recently, highly expressive classes of variational wave functions based on machine learning techniques—commonly referred to as neural-network quantum states (NNQS)—have been introduced [17]. While these ansätze can efficiently capture complex correlations [26, 27, 28, 29] and have been successfully used to simulate the quantum many-body dynamics of several systems using tVMC [17, 30, 31, 32, 33, 34, 35, 36], their application in this context often suffers from numerical instabilities [37]. It is well known that this problem can be mitigated by the use of regularization techniques [37, 38], which, although effective in stabilizing the simulation, introduce bias that is difficult to quantify.
@@
To address this issue, we build on the concept of the local-in-time error (LITE), a quantity recently introduced in the context of variational dynamics [39, 40], which can be evaluated within the tVMC framework and provides a local estimate of the accuracy of the variational evolution. The LITE provides an intrinsic estimate of the local accuracy of the time evolution at each simulation step. We employ this idea to develop an adaptive scheme—adaptive time-dependent variational Monte Carlo (atVMC)—which dynamically adjusts the number of variational parameters evolved during the tVMC simulation, based on their estimated relevance. Specifically, we derive an analytical criterion from the LITE to assess the importance of each parameter during the evolution, enabling us to retain only those contributing significantly to the dynamics while keeping the squared LITE below a prescribed threshold. We benchmark the performance of our method on the one-dimensional transverse-field Ising (TFI) model, demonstrating that accurate simulations can be achieved while evolving only a subset of the variational parameters at each timestep. Moreover, we show that our adaptive strategy reduces the need for aggressive regularization, thereby improving overall stability and accuracy.
@@ II Methods
In this section, we describe the theoretical and algorithmic framework developed in this work. We begin by reviewing the time-dependent variational Monte Carlo (tVMC) method, which provides the main foundation of our approach. We then introduce the local-in-time error (LITE), a quantity that measures the local accuracy of the variational time evolution and plays a central role in our adaptive scheme. Building upon these tools, we present the adaptive tVMC (atVMC) algorithm, which dynamically selects the subset of variational parameters to be updated at each time step of the tVMC simulation. The adaptive strategy aims to minimize the computational cost while maintaining a controlled accuracy, as quantified by the LITE. Specific procedures for adding or removing parameters from the evolution are discussed in detail, together with a criterion to assess the significance of each parameter, which allows us to suppress directions in parameter space that do not contribute meaningfully to the physical dynamics and may lead to instabilities.
@@ II.1 The time-dependent variational Monte Carlo method
The tVMC method employs a variational ansatz to represent the state of a quantum system, reducing the problem of describing its dynamics to determining the evolution of its time-dependent variational parameters [21, 19, 20]. We denote by $\left|\psi(\boldsymbol{\alpha}_{t})\right\rangle$ the variational ansatz for the quantum state and by $\psi(\boldsymbol{x};\boldsymbol{\alpha}_{t})=\langle\boldsymbol{x}|\psi(\boldsymbol{\alpha}_{t})\rangle$ the variational wave function, where $\boldsymbol{x}$ is a generic configuration vector and $\boldsymbol{\alpha}_{t}$ is a vector of time-dependent variational parameters, $\alpha_{k}=\alpha_{k}(t)\in\mathbb{C}$. The expectation value at time $t$ of any observable $\hat{O}$ can be obtained as
@@
The evolution of the parameters at time $t$ is determined by demanding that their time derivatives $\dot{\boldsymbol{\alpha}}_{t}$ minimize, to leading order in $dt$, the Fubini-Study distance [17]
@@
between the variationally-evolved state $\left|\psi(\boldsymbol{\alpha}_{t}+\dot{\boldsymbol{\alpha}}_{t}\,dt)\right\rangle$ and the local exact evolution $e^{-\frac{i}{\hbar}\hat{H}\,dt}\left|\psi(\boldsymbol{\alpha}_{t})\right\rangle$. This approach is equivalent to applying McLachlan’s variational principle [41, 42] and yields the following set of equations of motion for the variational parameters:
$i\hbar\sum_{k^{\prime}}S_{k,k^{\prime}}\dot{\alpha}_{k}=F_{k}\text{,}$ (3)
[…]
@@
It is well known, however, that the numerical solution of the linear system (3) may lead to jump-like instabilities in the simulation. These instabilities are attributed to the amplification of the stochastic noise affecting the Monte Carlo averages. Such amplification occurs because the equations of motion (3) are typically ill-conditioned, due to redundancy in the parametrization of the variational quantum state—a feature that is common in more expressive and capable ansätze that employ a large number of parameters [37]. Typically, such instabilities are mitigated through the application of regularization techniques [37, 38, 43]. However, any form of regularization inevitably introduces some degree of bias into the dynamics, making it desirable to limit the strength of regularization in order to maintain stability with minimal distortion.
We remark that the problem of instabilities can be circumvented by employing a fundamentally different method, the projected time-dependent variational Monte Carlo (ptVMC) [44, 45, 46, 47], which avoids the ill-conditioned linear system altogether. However, this approach may entail a considerably higher computational cost [48]. More recently, alternative strategies have been proposed that depart from conventional time-stepping schemes, relying instead on global-in-time variational formulations that optimize the entire quantum trajectory over a finite time interval, by minimizing a suitable cost function that enforces the Schrödinger dynamics [49, 50]. While these methods may offer improved stability and accuracy, they also require solving high-dimensional optimization problems, which may limit their applicability to long-time simulations or large systems.
@@ II.2 The local-in-time error
To develop our adaptive algorithm, we need a quantifiable measure of the error in a given variational approximation of the exact dynamics. A natural choice is the Fubini-Study distance from Eq. (2). We define the instantaneous error per unit time as
@@
This is a nonnegative quantity that characterizes the rate of departure between the locally exact solution and a state that evolves on the variational manifold. A somewhat lengthy but straightforward calculation (see the Supplemental material of Ref. [38]) shows that $\varepsilon_{\text{FS}}^{2}$ can be expressed in terms of the energy variance $\operatorname{Var}(\hat{H})$, the quantum geometric tensor $S_{k,k^{\prime}}$, the force vector $F_{k}$, and the time derivatives of the variational parameters $\dot{\boldsymbol{\alpha}}_{k}$. The result is
$\hfill\varepsilon_{\text{FS}}^{2}(t)$ $=\frac{1}{\hbar^{2}}\operatorname{Var}(\hat{H})+$ (5)
@@
Following the terminology of Refs. [39, 40], we call $\varepsilon(t)$ the local-in-time error (LITE). Since $S$ is positive semi-definite (being a covariance matrix), and $\varepsilon^{2}$ is nonnegative by definition, from Eq. (6) it follows that $0\leq\varepsilon^{2}(t)\leq\text{Var}(\hat{H})/\hbar^{2}$. Beyond serving as a local error diagnostic, the LITE also provides an a posteriori upper bound to the global error [51]
[…]
@@ II.3 Adaptive tVMC
The algorithm we introduce in this work enables the on-the-fly adjustment of the number of variational parameters that are actively evolved during a tVMC simulation. A frozen, or deactivated, parameter remains constant during the time evolution, simplifying the equations of motion (3) with the aim of improving their conditioning and, consequently, the numerical stability of the simulation. Obviously, changing the ansatz expressivity affects the accuracy of the variational description of the evolution of the quantum state: freezing a parameter will in general increase the LITE. Our goal is to use the smallest number of active parameters while ensuring that $\varepsilon^{2}(t)$, the LITE squared, remains below a predefined error threshold $\lambda^{2}_{\text{LITE}}$. The decision to freeze or unfreeze parameters is based on their level of importance, defined in terms of their impact on the LITE. Specifically, we estimate the importance of a parameter by how much its presence reduces the error, thus allowing for a time-dependent ranking of the parameters based on their relevance. The rationale behind this approach is twofold. First, not evolving less important parameters reduces the risk of instabilities during the numerical treatment of their evolution. Second, as a beneficial by-product, reducing the number of parameters to be updated reduces the computational complexity of the method. In the following, we detail how it is possible to accurately control, via the LITE estimation, the freezing or unfreezing of one or more parameters in the context of the tVMC method.
@@ II.3.1 Freezing one parameter
At each time step we compare the squared LITE, $\varepsilon^{2}(t)$, with the threshold, $\lambda^{2}_{\text{LITE}}$. The possibility of freezing a parameter is considered only if $\varepsilon^{2}(t)<\lambda^{2}_{\text{LITE}}$. In this case, we need to estimate the importance of each active parameter $\alpha_{j}$. For each $\alpha_{j}$ we perform a rearrangement of the parameter ordering so that $\alpha_{j}$ appears last. In the corresponding reordered basis, the $S$ matrix and the $F$ vector are expressed in block-matrix form as
@@
where we introduced the submatrix $\tilde{K}$, the diagonal element $[S^{-1}]_{j,j}$, and the off-diagonal vector $\boldsymbol{W}_{j}$.
@@
To compute $\tilde{S}^{-1}$ efficiently, we exploit the fact that the full inverse $S^{-1}$ has already been computed. Applying a standard inversion formula for block matrices [52], we obtain

### v5

@@ Abstract
We introduce an extension of the time-dependent variational Monte Carlo (tVMC) method that adaptively controls the expressivity of the variational quantum state during the simulation of the dynamics. This adaptive tVMC (atVMC) approach is specifically designed to enhance numerical stability when overparameterized variational ansätze lead to ill-conditioned equations of motion. Building on the concept of the local-in-time error (LITE)—a measure of the deviation between variational and exact evolution—we introduce a procedure to quantify each parameter’s contribution to reducing the LITE, using only quantities already computed in standard tVMC simulations. These relevance estimates guide the selective evolution of only the most significant parameters at each time step, while maintaining a prescribed level of accuracy. We benchmark the algorithm on quantum quenches in the one-dimensional transverse-field Ising model using both spin-Jastrow and restricted Boltzmann machine wave functions, with an emphasis on overparameterized regimes. The adaptive scheme significantly improves numerical stability and reduces the need for strong regularization, enabling reliable simulations with highly expressive variational ansätze.
@@ I Introduction
Understanding the dynamics of strongly correlated quantum many-body systems remains a fundamental and long-standing challenge in physics. In recent years, this problem has attracted renewed interest, driven by theoretical and experimental advances in the study of out-of-equilibrium phenomena [RevModPhys.83.863, PhysRevLett.105.015702, Heyl_2018, Zhang_time_crystal_2017, Choi_time_crystal_2017, Langen_2016, Abanin_2017, Bernien_Quantum_scars_2017], as well as by the rapid development of quantum technologies [RevModPhys.86.153, Preskill2018quantumcomputingin, ezratty2024understandingquantumtechnologies2024]. In particular, accurately describing quantum dynamics is essential to understanding the propagation of entanglement and coherence, and ultimately, to achieving control over them [entanglement_review]. Exact solutions to the time evolution of quantum many-body systems are limited to a few analytically tractable or numerically trivial cases. The primary obstacle is the exponential growth of the Hilbert space with the number of degrees of freedom, which makes a full description of the wave function computationally intractable. Consequently, a variety of numerical methods have been developed to circumvent this limitation, including tensor network techniques, quantum Monte Carlo, and variational approaches based on neural-network quantum states, all of which aim to efficiently parameterize the wave function in regimes where exact methods become intractable [SCHOLLWOCK201196_DMRG, orus2014practical, Orus2019, BeccaSorellaQuantumMC, Carleo_Solving_many-body_with_neural_networks, glasser2018neural].
@@
In this work, we focus on the time-dependent variational Monte Carlo (tVMC) method [Carleo_Becca_Schiró_Fabrizio_2012, Carleo_Becca_Sanchez_2014, Carleo_PhD_Thesis], which approximates the system’s wave function using a variational ansatz and, as described below, evolves it according to a variational principle. A key feature of tVMC is its flexibility in treating systems of arbitrary dimensionality, as its formulation does not rely on a specific geometry and connectivity. Moreover, since the wave function is represented by a parameterized variational ansatz, the computational cost of the method depends on the number of variational parameters rather than on the dimension of the Hilbert space, allowing for favorable scaling even in large or high-dimensional systems. This makes it complementary to other powerful methods, such as tensor networks [DMRG_PhysRevLett.69.2863, SCHOLLWOCK201196_DMRG, orus2014practical, Orus2019, PAECKEL2019167998, 10.21468/SciPostPhysLectNotes.8, PhysRevLett.126.170603], which are most effective in low-dimensional systems or when the entanglement is sufficiently constrained. However, the accuracy of a tVMC simulation crucially depends on the expressivity of the chosen variational ansatz. Recently, highly expressive classes of variational wave functions based on machine learning techniques—commonly referred to as neural-network quantum states (NNQS)—have been introduced [Carleo_Solving_many-body_with_neural_networks]. While these ansätze can efficiently capture complex correlations [PhysRevB.106.205136, Ureña_Sojo_Bermejo-Vega_Manzano_2024, Dash_Gravina_Vicentini_Ferrero_Georges_2025, PhysRevX.7.021021] and have been successfully used to simulate the quantum many-body dynamics of several systems using tVMC [Carleo_Solving_many-body_with_neural_networks, PhysRevB.103.165103, doi:10.1126/sciadv.abl6850, PhysRevLett.131.046501, PRXQuantum.4.040302, PhysRevB.110.104411, PhysRevB.109.235120, PhysRevB.109.155128], their application in this context often suffers from numerical instabilities [Hofmann_Role_of_stochastic_noise]. It is well known that this problem can be mitigated by the use of regularization techniques [Hofmann_Role_of_stochastic_noise, Schmitt_and_Heyl], which, although effective in stabilizing the simulation, introduce biases that are difficult to quantify. Addressing these issues is the main motivation of the present work.
@@
To this end, we build on the concept of the local-in-time error (LITE), a quantity recently introduced in the context of variational dynamics [Martinazzo_LITE, martinazzo2021commentregularizingmctdhequations], which can be evaluated within the tVMC framework and provides a local estimate of the accuracy of the variational evolution. The LITE provides an intrinsic estimate of the local accuracy of the time evolution at each simulation step. We employ this idea to develop an adaptive scheme—adaptive time-dependent variational Monte Carlo (atVMC)—which dynamically adjusts the number of variational parameters evolved during the tVMC simulation, based on their estimated relevance. Specifically, we derive an analytical criterion from the LITE to assess the importance of each parameter during the evolution. This enables us to retain only those parameters contributing significantly to the dynamics, while keeping the squared LITE below a prescribed threshold, thus effectively and dynamically filtering out redundant expressivity in the variational manifold. We benchmark the performance of our method on the one-dimensional transverse-field Ising (TFI) model, demonstrating that accurate simulations can be achieved while evolving only a subset of the variational parameters at each timestep. Moreover, we show that our adaptive strategy reduces the need for aggressive regularization, thereby improving overall stability and accuracy.
@@ II Methods
In this section, we describe the theoretical and algorithmic framework developed in this work. We begin by reviewing the time-dependent variational Monte Carlo (tVMC) method, which provides the main foundation of our approach. We then introduce the local-in-time error (LITE), a quantity that measures the local accuracy of the variational time evolution and plays a central role in our adaptive scheme. Building upon these tools, we present the adaptive tVMC (atVMC) algorithm, which dynamically selects the subset of variational parameters to be updated at each time step of the tVMC simulation. The adaptive strategy aims to reduce the number of variational parameters that can evolve in time while maintaining a controlled accuracy, as quantified by the LITE. Specific procedures for adding or removing parameters from the evolution are discussed in detail, together with a criterion to assess the significance of each parameter, which allows us to suppress directions in parameter space that do not contribute meaningfully to the physical dynamics and may lead to instabilities.
@@ II.1 The time-dependent variational Monte Carlo method
The tVMC method employs a variational ansatz to represent the state of a quantum system, reducing the problem of describing its dynamics to determining the evolution of its time-dependent variational parameters [Carleo_PhD_Thesis, Carleo_Becca_Schiró_Fabrizio_2012, Carleo_Becca_Sanchez_2014]. We denote by $\left|\psi(\boldsymbol{\alpha}_{t})\right\rangle$ the variational ansatz for the quantum state and by $\psi(\boldsymbol{x};\boldsymbol{\alpha}_{t})=\langle\boldsymbol{x}|\psi(\boldsymbol{\alpha}_{t})\rangle$ the variational wave function, where $\boldsymbol{x}$ is a generic configuration vector and $\boldsymbol{\alpha}_{t}$ is a vector of time-dependent variational parameters, $\alpha_{k}=\alpha_{k}(t)\in\mathbb{C}$. The expectation value at time $t$ of any observable $\hat{O}$ can be obtained as
@@
The evolution of the parameters at time $t$ is determined by demanding that their time derivatives $\dot{\boldsymbol{\alpha}}_{t}$ minimize, to leading order in $dt$, the Fubini–Study distance [Carleo_Solving_many-body_with_neural_networks, Provost1980]
@@
between the variationally-evolved state $\left|\psi(\boldsymbol{\alpha}_{t}+\dot{\boldsymbol{\alpha}}_{t}\,dt)\right\rangle$ and the local exact evolution $e^{-\frac{i}{\hbar}\hat{H}\,dt}\left|\psi(\boldsymbol{\alpha}_{t})\right\rangle$. This approach is equivalent to applying McLachlan’s variational principle [McLachlan, Yuan2019theoryofvariational] and yields the following set of equations of motion for the variational parameters:
$i\hbar\sum_{k^{\prime}}S_{k,k^{\prime}}\dot{\alpha}_{k^{\prime}}=F_{k}\text{,}$ (3)
where $S_{k,k^{\prime}}=\langle\mathcal{O}_{k}^{*}\mathcal{O}_{k^{\prime}}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{O}_{k^{\prime}}\rangle$ is the quantum geometric tensor and $F_{k}=\langle\mathcal{O}_{k}^{*}\mathcal{E}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{E}\rangle$ is the force vector. These are defined in terms of the local operators $\mathcal{O}_{k}(\boldsymbol{x};\boldsymbol{\alpha}_{t})=\frac{\partial}{\partial\alpha_{k}}\ln\langle\boldsymbol{x}|\psi(\boldsymbol{\alpha}_{t})\rangle$ and the local energy $\mathcal{E}(\boldsymbol{x};\boldsymbol{\alpha}_{t})=\langle\boldsymbol{x}|\hat{H}|\psi(\boldsymbol{\alpha}_{t})\rangle/\langle\boldsymbol{x}|\psi(\boldsymbol{\alpha}_{t})\rangle$. Expectation values, denoted by $\langle\cdot\rangle$, are computed with respect to the probability $p(\boldsymbol{x};\boldsymbol{\alpha}_{t})=|\psi(\boldsymbol{x};\boldsymbol{\alpha}_{t})|^{2}/\|\psi(\boldsymbol{\alpha}_{t})\|^{2}$ introduced above.
@@
It is well known, however, that the numerical solution of the linear system (3) may lead to jump-like instabilities in the simulation. These instabilities are attributed to the amplification of the stochastic noise affecting the Monte Carlo averages. Such amplification occurs because the equations of motion (3) are typically ill-conditioned, due to redundancy in the parameterization of the variational quantum state—a feature that is common in more expressive and capable ansätze that employ a large number of parameters [Hofmann_Role_of_stochastic_noise]. Typically, such instabilities are mitigated through the application of regularization techniques [Hofmann_Role_of_stochastic_noise, Schmitt_and_Heyl, quench_critical_point]. However, any form of regularization inevitably introduces some degree of bias into the dynamics, making it desirable to limit the strength of regularization in order to maintain stability with minimal distortion.
We remark that the problem of instabilities can be circumvented by employing a fundamentally different method, the projected time-dependent variational Monte Carlo (ptVMC) [Sinibaldi2023unbiasingtime, Gutierrez2022realtimeevolution, PhysRevA.108.022210, Nys_Pescia_Sinibaldi_Carleo_2024], which avoids the ill-conditioned linear system altogether. However, this approach may entail a considerably higher computational cost [Gravina2025neuralprojected]. More recently, alternative strategies have been proposed that depart from conventional time-stepping schemes, relying instead on global-in-time variational formulations that optimize the entire quantum trajectory over a finite time interval, by minimizing a suitable cost function that enforces the Schrödinger dynamics [Sinibaldi2024time, VandeWalle_2025]. While these methods may offer improved stability and accuracy, they also require solving high-dimensional optimization problems, which may limit their applicability to long-time simulations or large systems.
@@ II.2 The local-in-time error
To develop our adaptive algorithm, we need a quantifiable measure of the error in a given variational approximation of the exact dynamics. A natural choice is the Fubini–Study distance from Eq. (2). We define the instantaneous error per unit time as
@@
[…]
@@
Following the terminology of Refs. [Martinazzo_LITE, martinazzo2021commentregularizingmctdhequations], we call $\varepsilon(t)$ the local-in-time error (LITE). Since $S$ is positive semi-definite (being a covariance matrix), and $\varepsilon^{2}$ is nonnegative by definition, from Eq. (6) it follows that $0\leq\varepsilon^{2}(t)\leq\text{Var}(\hat{H})/\hbar^{2}$. Beyond serving as a local error diagnostic, the LITE also provides an a posteriori upper bound to the global error [Lubich_2008]
$\|\left|\psi(\boldsymbol{\alpha}_{t})\right\rangle-\left|\psi_{\text{ex}}(t)\right\rangle\|\leq\int_{0}^{t}\varepsilon(t^{\prime})\,dt^{\prime}\text{,}$ (7)
@@ II.3 Adaptive tVMC
The algorithm we introduce in this work enables the on-the-fly adjustment of the number of variational parameters that are actively evolved during a tVMC simulation. A frozen, or deactivated, parameter remains constant during the time evolution, simplifying the equations of motion (3) with the aim of improving their conditioning and, consequently, the numerical stability of the simulation. Obviously, changing the ansatz expressivity affects the accuracy of the variational description of the evolution of the quantum state: freezing a parameter will in general increase the LITE. Our goal is to use the smallest number of active parameters while ensuring that $\varepsilon^{2}(t)$, the LITE squared, remains below a predefined error threshold $\lambda^{2}_{\text{LITE}}$. The decision to freeze or unfreeze parameters is based on their level of importance, defined in terms of their impact on the LITE. Specifically, we estimate the importance of a parameter by how much its presence reduces the error, thus allowing for a time-dependent ranking of the parameters based on their relevance. The rationale behind this approach is twofold. First, not evolving less important parameters reduces the risk of instabilities during the numerical treatment of their evolution. Second, as a beneficial by-product, reducing the number of parameters to be updated reduces the computational cost of solving Eq.(3). In the following, we detail how it is possible to accurately control, via the LITE estimation, the freezing or unfreezing of one or more parameters in the context of the tVMC method.
@@ II.3.1 Freezing one parameter
At each time step, we compare the squared LITE, $\varepsilon^{2}(t)$, with the threshold, $\lambda^{2}_{\text{LITE}}$. The possibility of freezing a parameter is considered only if $\varepsilon^{2}(t)<\lambda^{2}_{\text{LITE}}$. In this case, we need to estimate the importance of each active parameter $\alpha_{j}$. For each $\alpha_{j}$ we perform a rearrangement of the parameter ordering so that $\alpha_{j}$ appears last. In the corresponding reordered basis, the $S$ matrix and the $F$ vector are expressed in block-matrix form as
@@
where we introduced the submatrix $\tilde{K}$, the diagonal element $S^{-1}]_{j,j}$, and the off-diagonal vector $\boldsymbol{W}_{j}$.
@@
To compute $\tilde{S}^{-1}$ efficiently, we exploit the fact that the full inverse $S^{-1}$ has already been computed. Applying a standard inversion formula for block matrices [Noble_linearAlgebra], we obtain